Two asteroids, one with a mass of 2.5x10^10 kg and another with a mass of 4x10^10 kg are seperated by 31km. What is the gravitational force the smaller mass exerts on the larger mass? Answer: 69.4 N
Answer #1

M1 = 2.5×10^10kg

M2 = 4×10^10kg

R = 31×10^3m

From F = G M1.M2/R^2 = 6.67×10^-11×2.5×10^10×4×10^10/(31000)^2

=) F = 69.4N